X+y=275
X*.1+y*.25=51.50

Help?

Answer by checkley79(3341)

X+Y=275 MULTIPLY BY -.1 & ADD.
.1X+.25Y=51.50
-.1X-.1Y=-27.50
------------------
.15Y=24
Y=24/.15
Y=160 ANS.
X+160=275
X=275-160
X=115 ANS.
PROOF:
.1*115+.25*160=51.50
11.50+40=51.50
51.50=51.50